There are 5 white socks and 3 black socks in a draw. Steven takes out 2 at random. Work out the probability that Steven takes out 2 socks of the same colour.

P(WW) = 5/8 x 4/7 = 20/56

P(BB) = 3/8 x 2/7 = 6/56

P(same colour) = 20/56 + 6/56 = 26/56 = 13/28 (simplified)

TJ

Related Maths GCSE answers

All answers ▸

Solve the simultaneous equations x^2 + y^2 = 9 and y = 3x + 3


Solve the quadratic equation (x^2)-7x+10=0


Solve the simultaneous equations x^2+y^2=1 and x+2y=1


Find the points where the curve given by: y = x^2 - 4x -12, and the line given by y = 2x - 12